OpenStudy (anonymous):

What expression gives the solutions of –5 + 2x² = –6x?

OpenStudy (anonymous):

rewrite is as: \[2x^2+6x-5=0\\ (2x+a)(x+b)=0\\ \text{where}\;a\times b=-10\qquad a+b=6\] or use the quadratic formula

OpenStudy (anonymous):

if\[ax^2+bx+c=0\\ \text{thetn}\quad x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

i dont think i quite understand

OpenStudy (anonymous):

1. rewrite the given equation in standard form.. \[-5+2x^2=-6x\\ 2x^2+6x-5=0\] ok till here?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then use the quadratic formula from above.

OpenStudy (anonymous):

a="2" b="6" and c="-5"

OpenStudy (anonymous):

you will obtain two values for "x"

OpenStudy (anonymous):

i lose it there. i dont understand quadratic equations at all